OpenStudy (anonymous):

Please help! Multiply and simplify: x^2+9/x+3 * x/x-3

OpenStudy (anonymous):

\[\frac{x^2+9}{x+3}\times \frac{x}{x-3}\]?

OpenStudy (anonymous):

if that is right, then just multiply because nothing factors and nothing cancels

OpenStudy (anonymous):

Woops, minus 9.

OpenStudy (anonymous):

how'd i guess?

OpenStudy (anonymous):

now it factors as \[\frac{(x+3)(x-3)x}{(x+3)(x-3)}=x\]

OpenStudy (anonymous):

Oh I see. Thank you :)

OpenStudy (anonymous):

yw